OpenStudy (anonymous):

Express each ratio as a fraction in simplest form. 12 : 36

OpenStudy (anonymous):

answers could be 1/3 2/6 2/3 3/4

hartnn (hartnn):

does 12 goes into 36 ? how many times ?

OpenStudy (anonymous):

3 times

hartnn (hartnn):

yes! so, 12:36 = 12/36 = 12/ (3* 12) what gets cancelled ? what remains ?

OpenStudy (anonymous):

im not sure could you explain a bit more?

hartnn (hartnn):

yes, we wrote 36 as 12*3 right ? so, numerator and denominator both have which number ?

OpenStudy (anonymous):

3! right?

hartnn (hartnn):

the common number is 12 which gets cancelled and what remains is 1/3 got this ?

OpenStudy (anonymous):

ok i think i got it thanks so much i have one more problem for you if you have the time ok ?

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

4/7 = 3+x/21

hartnn (hartnn):

3+x/21 or (3+x)/21 ??

OpenStudy (anonymous):

x+3/21

OpenStudy (anonymous):

answers could be x = 9 x = 15 x=33 and 3/4 x=91

hartnn (hartnn):

did you try ?

OpenStudy (anonymous):

i am now what are your thoughts on this one?

OpenStudy (anonymous):

i think its x =9

OpenStudy (anonymous):

that would = the same as 4/7 as a decimal do you concur?

hartnn (hartnn):

if it is 4/7 = (3+x)/21 then x =9

hartnn (hartnn):

and i think its 4/7 = (3+x)/21 only

OpenStudy (anonymous):

ok what if it does not have the ( )

hartnn (hartnn):

then answer is complicated negative value which is not in choices..

hartnn (hartnn):

4/7 -3 = x/21 (4-21)*3 = x ....not in choices...

OpenStudy (anonymous):

ok x=9 then

hartnn (hartnn):

yes

OpenStudy (anonymous):

Ruiz invests $550 at a simple interest rate of 4%. How much interest will he earn after 5 years? will you help with this last one too?

OpenStudy (anonymous):

answers could be $11 $110 $1100 $11,000